A 24% trade discount is equal to $365. What is the list price (rounded to the nearest dollar)?

a. $453       b. $1,156         c. $993            d. $1,799         e. $1,521

 

 

7.      An invoice for $5,000 dated July 25, terms 5/10, net 30 is paid on August 4. What amount must be paid to reduce the debt to $1,000?

a. $1,000     b. $3,800         c. $4,995         d. $4,998         e. $5,000

 

 

8.      How much interest is owed on a loan of $11,500 borrowed at 5.5% from March 1, 2020 to August 15, 2020?

a. $11,683.68     b. $318.86      c. $212.34      d. $183.68       e. None of these

         

 

 

 

 

9.      How much must be placed in a two-month term deposit earning 4.8% simple interest in order to earn $275 interest?

a. $12,500        b. $5,729.17         c. $23,333.33        d. $34,375.00      e. $53,333.33

 

 

10.  In how many months will $2,500 earn $51.04 interest at 3.5% simple interest?

a. 5             b. 7           c. 9          d. 6            e. None of these

 

1
Expert's answer
2021-02-24T11:24:24-0500

6.

"\frac{365}{0.24}=1521 ~"$ (e).

7.

"(5000-1000)\cdot 0.95=3800~" $ (b).

8.

"11500\cdot 0.055 \cdot \frac{168}{366}=290.33" $ (e).

9.

"\frac{275}{0.048}\cdot \frac{12}{2}=34375" $ (d).

10.

"\frac{51.04}{2500}\cdot \frac{100}{3.5}\cdot 12=7" (b).